OpenStudy (maheshmeghwal9):

Calculate the wavelength of the spectral line obtained in the spectrum of \[Li^{2+}\] ion when the transition takes place between two levels whose sum is 4 & the difference is 2.

OpenStudy (maheshmeghwal9):

I know this formula: -\[\frac{1}{\lambda}=R \times Z^2 \times [\frac{1}{n_1^2}-\frac{1}{n_2^2}].\]Where \[n_2>n_1.\]

OpenStudy (maheshmeghwal9):

Therefore I took the question as \[n_2+n_1=4\]\[n_2-n_1=2.\]

OpenStudy (maheshmeghwal9):

I gt \[n_2=3; \space \space n_1=1.\]

OpenStudy (maheshmeghwal9):

but when I put these values in the formulae I didn't gt answer.

OpenStudy (maheshmeghwal9):

I know this too: - \[Li^{2+}.\] \[Z=3.\]

OpenStudy (maheshmeghwal9):

So anybody plz help me:)

OpenStudy (maheshmeghwal9):

@mukushla @mathslover @vishweshshrimali5 @PaxPolaris @UnkleRhaukus @A.Avinash_Goutham @JFraser Please help:)

OpenStudy (paxpolaris):

I didn't even know lithium mad a 2+ ion ... :S

OpenStudy (maheshmeghwal9):

ohhhh

OpenStudy (vishweshshrimali5):

U can atleast find the transition i.e. lower and upper level i mean initial and final energy level by simple maths

OpenStudy (vishweshshrimali5):

what do you get the value of lower and highest energy level for transition

OpenStudy (maheshmeghwal9):

I have given that above:)

OpenStudy (vishweshshrimali5):

sorry didn't read that

OpenStudy (maheshmeghwal9):

np:)

OpenStudy (vishweshshrimali5):

u have done everything correctly ur answer should be correct

OpenStudy (maheshmeghwal9):

oh ok but would u also give me the answer plz? so i can match mine.

OpenStudy (vishweshshrimali5):

I would have to calculate wait

OpenStudy (maheshmeghwal9):

ok:)

OpenStudy (vishweshshrimali5):

what did u take the value of R = ?

OpenStudy (maheshmeghwal9):

R=109667 cm^-1

OpenStudy (paxpolaris):

i got 11.4 nm http://lmgtfy.com/?q=1%2F8%2F(Rydberg+Constant)

OpenStudy (maheshmeghwal9):

But answer given is 1.14 x 10^-6 cm.

OpenStudy (paxpolaris):

that's the same...

OpenStudy (maheshmeghwal9):

oh ok

OpenStudy (maheshmeghwal9):

But how?

OpenStudy (paxpolaris):

http://lmgtfy.com/?q=1%2F8%2F(Rydberg+Constant)+in+cm

OpenStudy (vishweshshrimali5):

yes \[1.14 * 10^{-6} cm = 1.14 * 10^{-8} m = 11.4 * 10 ^{-9} m = 11.4 nm\]

OpenStudy (ujjwal):

Its 11.4 nm Put R=10973731.57 and that's it.

OpenStudy (maheshmeghwal9):

But R = 109667 cm^{-1}

OpenStudy (vishweshshrimali5):

http://bit.ly/NcPY8j

OpenStudy (paxpolaris):

there are 10^7 nanometers in 1 centimeter 11.4nm = 1.14x10^-6 cm

OpenStudy (maheshmeghwal9):

oh ya i gt my answer by R =109677 cm^{-1} Thanx to all :D

OpenStudy (vishweshshrimali5):

:D

OpenStudy (paxpolaris):

not that it matter for this question ... I took R=109737 cm^{-1} http://lmgtfy.com/?q=(Rydberg+Constant)+%20in+cm%5E-1